The annual premium for a $5,000 insurance policy against the theft of a painting is $200. If the (empirical) probability that the painting will be stolen during the year is 0.02, what is your expected return from the insurance company if you take out this insurance? Let X be the random variable for the amount of money received from the insurance company in the given year. Edollars
Answer #1

Ans:

Probability that painting will be stolen=0.02

Probability that painting will not be stolen=1-0.02=0.98

If painting is stolen,then amount paid by insurance company is =5000

If painting is not stolen,then amount received from insurance company=-200

Expected return=0.02*5000-0.98*200

=100-196

=-96